I'm in 9th grade.
My math teacher said that there is another way to ADD/Subtract improper fractions without making them proper by adding/subtracting the whole numbers and then the fractions. But on my Tests, whenever I do this, I get the answer incorrect.I'm not sure if I'm using this method incorrectly, we never really took any notes on it. Can Anyone tell me how to do this?

Found 2 solutions by Mathtut, stanbon:
Answer by Mathtut(3670)

This is fairly straight forward if the fraction you are subtracting is smaller than the fraction your subtract it from
:
let say 12 1/2- 9 1/8
:
that would be 3 when subtracting the whole numbers. for the fractions 1/2-1/8, we need a LCD which is 8 in this case. 4/8-1/8=3/8 ...so you have 3 3/8 for an answer
:
but what if the fraction being subtracted is greater than the fraction from which it being subtracted from
:
such as:
:
29 1/3 - 22 3/8
:
you can subtract 22 from 29 and get 7
:
then you have 1/3-3/8...you must make them like terms with a common denominator
:
24 would be the LCD
:
8/24-9/24=-1/24
:
which means you have to subtract that from 7 and that makes 6 23/24
:
for adding
:
adding the fraction parts that are less than one is straightforward but if you add two fractions that end up greater than one then you have to add one to your whole answer and then add what is left over to the new whole number.
:
such as
:
10 3/8+ 12 7/8
:
22 for the whole part 3/8+7/8=10/8. subtract one which is 8/8 from this and add it to your whole answer of 22+1=23 and you have 2/8=1/4 left over
:
so 23 1/4

Answer by stanbon(75887)

My math teacher said that there is another way to ADD/Subtract improper fractions without making them proper by adding/subtracting the whole numbers and then the fractions. But on my Tests, whenever I do this, I get the answer incorrect.I'm not sure if I'm using this method incorrectly, we never really took any notes on it. Can Anyone tell me how to do this?
------------------------------
Write the problem vertically (one below the other):
Here is a SUBTRACTION Example:
..........
2 3/5
-1 1/2
--------------
Rewrite the problem with the fractions having the same denominator:
2 (6/10)
-1 (5/10)
----------------
1 1/10
=========================
Cheers,
Stan H.